# Establishing Linear Surrogate Regret Bounds for Convex Smooth Losses via Convolutional Fenchel–Young Losses

### v2

@@ 2.2 Surrogate Losses and Surrogate Regret Bounds
$\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\leq\psi(\mathtt{Regret}_{L}(\bm{\theta},\bm{\eta})),~{}~{}~{}~{}~{}\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K}.$ (6)
While a linear regret rate $\psi(r)=\mathcal{O}(r)$ is the best possible, previously known examples are either non-smooth or non-convex surrogate losses, including the (non-smooth) hinge loss [6] and (non-convex) sigmoid loss [15]. These loss functions may face challenges in optimization and estimation [63]. In this work, we aim to develop a framework that facilitates the use of convex smooth surrogates with linear surrogate regret bounds.
@@ 2.3 Fenchel–Young Loss
In this work, we build upon Fenchel–Young losses [10] to construct convex smooth surrogates equipped with linear regret rate functions. Let us review Fenchel–Young losses first.
@@
The function $\Omega$ is often regarded as a negentropy, which equals the Bayes risk of its induced Fenchel–Young loss. Fenchel–Young losses often impose additional requirements on the domains of $\Omega$ and $\Omega^{*}$, tailored to specific target problems, and we will highlight them when necessary.
Fenchel–Young losses possess favorable properties: they are inherently convex (in score $\bm{\theta}$) by definition. Moreover, they can induce Fisher-consistent estimators of $\mathbb{E}_{Y\sim\bm{\eta}}[\bm{\rho}(Y)]$, which recovers the class distribution $\bm{\eta}$ under multiclass classification with $\bm{\rho}(y)=\bm{e}_{y}$; or the linear property in general [1]. Many common losses are encompassed in Fenchel–Young losses, including the cross-entropy loss, squared loss, and Crammer–Singer loss [20]. For example, $\Omega(\bm{p})=\frac{1}{2}\|\bm{p}\|^{2}+\mathbb{I}_{\Delta^{K}}(\bm{p})$ generates the sparsemax loss, which induces the sparsemax as a Fisher-consistent probability estimator of $\bm{\eta}$ [41].
@@ 3 Convex Smooth Surrogates with Linear Surrogate Regret Bounds
In this section, we first show our construction of surrogate loss and prediction links. After showing that this loss is convex and smooth with an appropriately chosen base negentropy (Theorem 11), we move on to the main result of this paper, linear surrogate regret bounds with convex smooth surrogate losses (Theorem 13). We further discuss the computational aspects and improve the regret bound constant. The missing proofs of technical results are deferred to Appendix A.
@@ 3.1 Construction
We design a convex smooth surrogate loss built upon the framework of Fenchel–Young losses. To make its surrogate regret bound linear, we craft an negentropy function by delicately leveraging the structure of a target prediction problem. Denote by $T$ the polyhedral convex function $T(\bm{p})=-\min_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$. Then, we have the following definition:
@@
Suppose $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition for a target loss $\ell$. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$, its convolutional negentropy is defined as follows:
@@
While commonly used negentropy functions in Fenchel–Young losses, such as the Shannon negentropy and norm negentropy [12], are unstructured toward a target loss, the convolutional negentropy $\Omega_{T}$ encodes a target loss $\ell$ explicitly into the base negentropy $\Omega$ via $T$. This polyhedral convex function $T$ is an affinely transformed Bayes risk of a target loss $\ell$. Indeed, with the linear property $\bm{p}=\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$, we have
@@
$\Omega$ is proper convex and lower-semicontinuous (l.s.c.), and satisfies $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))\subseteq\mathrm{dom}(\Omega)$ and $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{\varrho}$.
It is a mild requirement. Indeed, proper convexity and lower-semicontinuity merely aim to avoid pathologies, and the domain assumptions are met by a differentiable and finite $\Omega$ over the valid prediction space $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))$. The domain assumption on $\Omega^{*}$ is crucial for the induced loss to have an unconstrained domain $\mathbb{R}^{\varrho}$. Then we show an explicit form of the conjugated convolutional negentropy.
@@
Suppose Condition 1 holds, then $\Omega_{T}^{*}$ is proper convex and l.s.c. with $\mathrm{dom}(\Omega_{T}^{*})=\mathbb{R}^{\varrho}$, and can be expressed as follows:
$\Omega_{T}^{*}(\bm{\theta})=\inf_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\quad\text{for any $\bm{\theta}\in\mathbb{R}^{\varrho}$},$ (10)
where $\mathcal{L}^{\bm{\rho}}\in\mathbb{R}^{\varrho\times N}$ is the loss matrix with the $t$-th column being $\bm{\ell}^{\bm{\rho}}(t)\in\mathbb{R}^{\varrho}$ for $t\in\widehat{\mathcal{Y}}$. In addition, there always exists $\bm{\pi}\in\Delta^{N}$ that achieves the infimum of (10) for any $\bm{\theta}\in\mathbb{R}^{\varrho}$.
The conjugated form (10) owes to the infimal convolution between the base negentropy $\Omega$ and the transformed target Bayes risk $\min_{t\in\widehat{\mathcal{Y}}}\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\rangle$. We coined the name of the convolutional negentropy inspired by this structure. Intuitively, the conjugated convolutional negentropy $\Omega_{T}^{*}$ can be viewed as a “perturbed” version of the conjugated base negentropy $\Omega^{*}$ toward the direction of the target loss matrix $\mathcal{L}^{\bm{\rho}}$. This result facilitates a more concrete formulation of the Fenchel–Young loss generated by $\Omega_{T}$, which we refer to as the convolutional Fenchel–Young loss—the central focus of this work.
@@
Suppose that a discrete target loss $\ell$ enjoys $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$ satisfying Condition 1, the surrogate loss $L_{\Omega_{T}}:\mathbb{R}^{\varrho}\times\mathcal{Y}\to\overline{\mathbb{R}}$ induced by the convolutional negentropy $\Omega_{T}$ is defined as follows:
@@
Given a surrogate loss, we need to specify a prediction link. For general discrete prediction problems where $\mathcal{Y}\neq\widehat{\mathcal{Y}}$, a prediction link yields a discrete prediction in $\widehat{\mathcal{Y}}$ from a score $\bm{\theta}\in\mathbb{R}^{\varrho}$ obtained through minimizing the surrogate loss. Thus a loss and link are the two sides of the same coin. Unlike the standard argmax link in multiclass classification [65, 68], we create an alternative argmax-like prediction link based on the minimizer of (10).
@@
Let $\Pi:\mathbb{R}^{\varrho}\to 2^{\Delta^{N}}$ be a set-valued map defined as follows:
$\Pi(\bm{\theta})\coloneqq\mathop{\mathrm{missing}}{argmin}_{\bm{\pi}\in\Delta^{N}}\,\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\quad\text{for any $\bm{\theta}\in\mathbb{R}^{\varrho}$}.$ (12)
Let $\bm{\pi}:\mathbb{R}^{\varrho}\to\Delta^{N}$ be a selector of $\Pi$ such that $\bm{\pi}(\bm{\theta})\in\Pi(\bm{\theta})$ for all $\bm{\theta}\in\mathbb{R}^{\varrho}$. Then the $\bm{\pi}$-argmax link $\varphi:\mathbb{R}^{\varrho}\to\widehat{\mathcal{Y}}$ is defined as
$\varphi(\bm{\theta})\in\mathop{\mathrm{missing}}{argmax}_{t\in\widehat{\mathcal{Y}}}\,\pi_{t}(\bm{\theta})\quad\text{for any $\bm{\theta}\in\mathbb{R}^{\varrho}$},$
@@
The existence of such links is guaranteed as long as $\Pi(\bm{\theta})$ is non-empty for every $\bm{\theta}\in\mathbb{R}^{\varrho}$, which is guaranteed by Lemma 8. While the standard argmax link returns a prediction $t\in\widehat{\mathcal{Y}}$ with the maximum score $\theta_{t}$, the ${\bm{\pi}}$-argmax link returns $t$ with the maximum $\pi_{t}$. This probabilistic quantity $\bm{\pi}\in\Pi(\bm{\theta})$ defined via (10) distorts the original score $\bm{\theta}$ by leveraging the target loss structure $\mathcal{L}^{\bm{\rho}}$.
@@ 3.2 Convexity and Smoothness
Before discussing surrogate regret bounds, we verify that convolutional Fenchel–Young losses are indeed convex and smooth, implied by the conjugacy between smoothness and strong convexity [28].
@@
Suppose that a discrete target loss $\ell$ enjoys $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For a base negentropy $\Omega$, we additionally suppose that Condition 1 is satisfied. If $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$, $L_{\Omega_{T}}(\cdot,y)$ is convex and differentiable over $\mathbb{R}^{\varrho}$ for any $y\in\mathcal{Y}$. If $\Omega$ is additionally strongly convex on $\mathrm{dom}(\Omega)$, $L_{\Omega_{T}}(\cdot,y)$ is smooth over $\mathbb{R}^{\varrho}$ for any $y\in\mathcal{Y}$.
There are several exemplar base entropies fulfilling the conditions of Theorem 11. For example, the squared norm $\Omega(\bm{p})=\frac{1}{2}\|\bm{p}\|^{2}$ is strongly convex with the self-conjugate $\Omega^{*}(\bm{\theta})=\frac{1}{2}\|\bm{\theta}\|^{2}$. In this case, $\Omega$ is effective over the entire $\mathbb{R}^{\varrho}$ and hence includes $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))$, satisfying Condition 1. In light of the target-loss decomposition (3), we have $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))=\Delta^{K}$. In this case, the Shannon negentropy $\Omega(\bm{p})=\langle\bm{p},\ln\bm{p}\rangle+\mathbb{I}_{\Delta^{K}}(\bm{p})$ is strongly convex on $\Delta^{K}$ with its conjugate $\Omega^{*}(\bm{\theta})=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle$ satisfying $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{K}$, where $\ln$ and $\exp$ are element-wise. Hence, both base negentropy functions yield convex and smooth $L_{\Omega_{T}}$.
While the surrogate loss $L_{\Omega_{T}}$ is now ensured to be convex and smooth, its gradient calculation, which is the basis for gradient-based optimization [11, 9], remains non-trivial. This is due to the gradient of the term $\min_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$, which cannot be written analytically in general. We show that its gradient calculation can be reduced to the calculation of $\Pi(\bm{\theta})$ through the following variant of envelope theorems.
@@
Suppose Condition 1 holds and $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$. For any $\bm{\theta}\in\mathbb{R}^{\varrho}$ and $\bm{\pi}\in\Pi(\bm{\theta})$, we have
@@
Deviating from the standard envelope theorems [9, 7], we carefully addresses the non-uniqueness of the minimizers. Given this, the gradient of the convolutional Fenchel–Young loss (11) is accessed via $\nabla_{\bm{\theta}}L_{\Omega_{T}}(\bm{\theta},y)=\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\bm{\rho}(y)$. At the core of its proof, the differentiability of $\Omega_{T}^{*}$ (indirectly obtained via Theorem 11) is vital.

### v3

@@ 2.2 Surrogate Losses and Surrogate Regret Bounds
$\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\leq\psi(\mathtt{Regret}_{L}(\bm{\theta},\bm{\eta})),~~~~~\text{for any~}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{d}\times\Delta^{K}.$ (6)
While a linear regret rate $\psi(r)=\mathcal{O}(r)$ is the best possible, previously know linear-rate losses are either non-smooth or non-convex, including the (non-smooth) hinge loss [06_calibration] and (non-convex) sigmoid loss [Nontawat]. These loss functions may face challenges in optimization and estimation [smooth1]. In this work, we aim to develop a framework that facilitates the use of convex smooth surrogates with linear surrogate regret bounds.
@@ 2.3 Fenchel–Young Loss
In this work, we build upon Fenchel–Young losses [FY, energy_fy, sparse_FY] to construct convex smooth surrogates equipped with linear regret rate functions. Let us review Fenchel–Young losses first.
@@
Similar constructions can also be found in duchi2018multiclass, least_square, which focus on multiclass classification. The function $\Omega$ is often regarded as a generalized negentropy, which often equals the negative of the Bayes risk of its induced Fenchel–Young loss. Fenchel–Young losses often impose additional requirements on the domains of $\Omega$ and $\Omega^{*}$, tailored to specific target problems, and we will highlight them when necessary.
Fenchel–Young losses possess favorable properties: they are inherently convex (in score $\bm{\theta}$) by definition. Moreover, we can use the map $\nabla\Omega^{*}(\bm{\theta})$ to obtain the mean $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$, which is the class distribution $\bm{\eta}$ under multiclass classification with $\bm{\rho}(y)=\bm{e}_{y}$; or linear properties in general [abernethy2012characterization]. Many common losses are encompassed in Fenchel–Young losses, including the cross-entropy loss, squared loss, and Crammer–Singer loss [crammer]. For example, $\Omega(\bm{p})=\frac{1}{2}\|\bm{p}\|^{2}+\mathbb{I}_{\Delta^{K}}(\bm{p})$ generates the sparsemax loss, which induces the sparsemax as a Fisher-consistent estimator of $\bm{\eta}$ [msparsemax].
Connections to information geometry.
While Fenchel–Young losses are systematically defined under the convex-analytic formulation, it also has a longstanding history in information geometry, particularly through its connection to the Bregman divergence and related generalizations [Fitzpatrick]. In FY, it is noted that a Fenchel–Young loss can be interpreted as the mixed-type Bregman divergence [it2]. When $\Omega$ is of Legendre-type, $\Omega$ can yield the dual coordinate system and the Fenchel–Young loss is further equivalent to the canonical divergence generated by $\Omega$ [IT1, Eq. (3.44)].
@@ 3 Convex Smooth Surrogates with Linear Surrogate Regret Bounds
In this section, we first show our construction of surrogate losses and prediction links. After showing that this loss is convex and smooth with an appropriately chosen base negentropy (Theorem 11), we move on to the main result of this paper, linear surrogate regret bounds with convex smooth surrogate losses (Theorem 13). We further discuss the computational aspects and improve the regret bound constant. The missing proofs are deferred to Appendix B.
@@ 3.1 Construction
We design a convex smooth surrogate loss built upon the framework of Fenchel–Young losses. To make its surrogate regret bound linear, we craft a negentropy by delicately leveraging the structure of a target prediction problem. Denote by $T$ the polyhedral convex function $T(\bm{p})=-\min_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$. Then, we have the following definition.
@@
Suppose $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition for a target loss $\ell$. For a negentropy $\Omega:\mathbb{R}^{d}\to\overline{\mathbb{R}}$, its convolutional negentropy is defined as follows:
@@
While commonly used negentropy in Fenchel–Young losses, such as the Shannon negentropy and norm negentropy [norm], are unstructured toward a target loss, the convolutional negentropy $\Omega_{T}$ encodes a target loss $\ell$ explicitly into the base negentropy $\Omega$ via $T$. This polyhedral convex function $T$ is an affinely transformed negative Bayes risk of a target loss $\ell$ when $\bm{p}\in\text{conv}\{\bm{\rho}(\mathcal{Y})\}$. Indeed, with the linear property $\bm{p}=\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$, i.e., expectation of $\bm{\rho}(y)$ w.r.t. class probability $\bm{\eta}$:
@@
$\Omega$ is proper convex and lower-semicontinuous (l.s.c.), and satisfies $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))\subseteq\mathrm{dom}(\Omega)$ and $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{d}$.
It is a mild requirement. Indeed, proper convexity and lower-semicontinuity merely aim to avoid pathologies, and the domain assumptions are met by a differentiable and finite $\Omega$ over the valid prediction space $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))$. The assumption on $\Omega^{*}$ is crucial for the induced loss to have unconstrained domain $\mathbb{R}^{d}$. Then we show an explicit form of the conjugated convolutional negentropy.
@@
Suppose Condition 1 holds, then $\Omega_{T}^{*}$ is proper convex and l.s.c. with $\mathrm{dom}(\Omega_{T}^{*})=\mathbb{R}^{d}$, and can be expressed as follows:
$\Omega_{T}^{*}(\bm{\theta})=\inf_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\quad\text{for any $\bm{\theta}\in\mathbb{R}^{d}$},$ (10)
where $\mathcal{L}^{\bm{\rho}}\in\mathbb{R}^{d\times N}$ is the loss matrix with the $t$-th column being $\bm{\ell}^{\bm{\rho}}(t)\in\mathbb{R}^{d}$ for $t\in\widehat{\mathcal{Y}}$. In addition, there always exists $\bm{\pi}\in\Delta^{N}$ that achieves the infimum of (10) for any $\bm{\theta}\in\mathbb{R}^{d}$.
The conjugated form (10) owes to the infimal convolution between the base negentropy $\Omega$ and the target Bayes risk $T$. We coined the name of the convolutional negentropy inspired by this structure. Intuitively, the conjugated convolutional negentropy $\Omega_{T}^{*}$ can be viewed as a “perturbed” version of the conjugated base negentropy $\Omega^{*}$ toward the direction of the target loss matrix $\mathcal{L}^{\bm{\rho}}$. This result facilitates a more concrete formulation of the Fenchel–Young loss generated by $\Omega_{T}$, which we refer to as the convolutional Fenchel–Young loss—the central focus of this work.
@@
Suppose that a discrete target loss $\ell$ enjoys $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For a negentropy $\Omega:\mathbb{R}^{d}\to\overline{\mathbb{R}}$ satisfying Condition 1, the convolutional Fenchel–Young loss $L_{\Omega_{T}}:\mathbb{R}^{d}\times\mathcal{Y}\to\overline{\mathbb{R}}$ induced by the convolutional negentropy $\Omega_{T}$ is defined as follows:
@@
Given a surrogate loss, we need to specify a prediction link. For general discrete prediction problems where $\mathcal{Y}\neq\widehat{\mathcal{Y}}$, a prediction link yields a discrete prediction in $\widehat{\mathcal{Y}}$ from a score $\bm{\theta}\in\mathbb{R}^{d}$ obtained through minimizing the surrogate loss. Thus, a loss and link are the two sides of the same coin. Unlike the standard argmax link in multiclass classification [07_multi, wang_unified], we create an alternative argmax-like prediction link based on the minimizer of (10).
@@
Let $\Pi:\mathbb{R}^{d}\to 2^{\Delta^{N}}$ be a set-valued map defined as follows:
$\Pi(\bm{\theta})\coloneqq\mathop{\mathrm{argmin}}\left\{\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}):\bm{\pi}\in\Delta^{N}\right\}\quad\text{for any $\bm{\theta}\in\mathbb{R}^{d}$}.$ (12)
Let $\bm{\pi}:\mathbb{R}^{d}\to\Delta^{N}$ be a selector of $\Pi$ such that $\bm{\pi}(\bm{\theta})\in\Pi(\bm{\theta})$ for all $\bm{\theta}\in\mathbb{R}^{d}$. [^4] We slightly abuse the notation $\bm{\pi}$ by using it for both a vector and a vector-valued mapping; in particular, the $\bm{\pi}$ appearing in the $\bm{\pi}$-argmax link refers to the selector function. Then the $\bm{\pi}$-argmax link $\varphi:\mathbb{R}^{d}\to\widehat{\mathcal{Y}}$ is defined as
$\varphi(\bm{\theta})\in\mathop{\mathrm{argmax}}\left\{\pi_{t}(\bm{\theta}):t\in\widehat{\mathcal{Y}}\right\}\quad\text{for any $\bm{\theta}\in\mathbb{R}^{d}$},$
@@
The existence of such links is guaranteed as long as $\Pi(\bm{\theta})$ is non-empty for every $\bm{\theta}\in\mathbb{R}^{d}$, which is guaranteed by Lemma 8. While the standard argmax link returns a prediction $t\in\widehat{\mathcal{Y}}$ with the maximum score $\theta_{t}$, the $\bm{\pi}$}-argmax link returns $t$ with the maximum $\pi_{t}$. This probabilistic quantity $\bm{\pi}\in\Pi(\bm{\theta})$ defined via (10) distorts the original score $\bm{\theta}$ by leveraging the target loss structure $\mathcal{L}^{\bm{\rho}}$.
@@ 3.2 Convexity and Smoothness
Before discussing surrogate regret bounds, we verify that convolutional Fenchel–Young losses are indeed convex and smooth, implied by the conjugacy between smoothness and strong convexity [dual-sm-cvx].
@@
Suppose that a discrete target loss $\ell$ enjoys $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For a base negentropy $\Omega$, we additionally suppose that Condition 1 is satisfied. If $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$, $L_{\Omega_{T}}(\cdot,y)$ is convex and differentiable over $\mathbb{R}^{d}$ for any $y\in\mathcal{Y}$. If $\Omega$ is additionally strongly convex on $\mathrm{dom}(\Omega)$, $L_{\Omega_{T}}(\cdot,y)$ is smooth over $\mathbb{R}^{d}$ for any $y\in\mathcal{Y}$.
There are several exemplar base negentropies fulfilling the conditions of Corollary 11. For example, the squared norm $\Omega(\bm{p})=\frac{1}{2}\|\bm{p}\|^{2}$ is strongly convex with the self-conjugate $\Omega^{*}(\bm{\theta})=\frac{1}{2}\|\bm{\theta}\|^{2}$. In this case, $\Omega$ is effective over the entire $\mathbb{R}^{d}$ and hence includes $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))$, satisfying Condition 1. In light of the target-loss decomposition (3), we have $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))=\Delta^{K}$. In this case, the Shannon negentropy $\Omega(\bm{p})=\langle\bm{p},\ln\bm{p}\rangle+\mathbb{I}_{\Delta^{K}}(\bm{p})$ is strongly convex on $\Delta^{K}$ with its conjugate $\Omega^{*}(\bm{\theta})=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle$ satisfying $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{K}$, where $\ln$ and $\exp$ are element-wise. Hence, both base negentropy yield convex and smooth $L_{\Omega_{T}}$.
Note that $L_{\Omega_{T}}$ is not locally strongly convex at every point. We can see this by noting that $T$ is not differentiable and neither is $\Omega_{T}$, which implies that $\Omega_{T}^{*}$ is not strictly convex [variational, Theorem 11.13]. This is important for establishing linear surrogate regret bounds (in Section 3.3) because the known square-root regret rate lower bound considers locally strongly convex surrogate losses [poly_regret, Theorem 2].
While the loss $L_{\Omega_{T}}$ is now ensured to be convex and smooth, its gradient calculation, which is the basis for gradient-based optimization [efficient, envelope], remains non-trivial. This is because the gradient of $\min_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ contained in (11) cannot be written analytically in general. We show that its gradient calculation reduces to computing $\Pi(\bm{\theta})$ through the following variant of envelope theorems.
@@
Suppose Condition 1 holds and $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$. For any $\bm{\theta}\in\mathbb{R}^{d}$ and $\bm{\pi}\in\Pi(\bm{\theta})$, we have
@@
Deviating from the standard envelope theorems [envelope, bertsekas], we carefully addresses the non-uniqueness of the minimizers. Given this, the gradient of the convolutional Fenchel–Young loss (11) is accessed via $\nabla_{\bm{\theta}}L_{\Omega_{T}}(\bm{\theta},y)=\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\bm{\rho}(y)$. At the core of its proof, the differentiability of $\Omega_{T}^{*}$ (indirectly obtained via Theorem 11) is vital.